What is the answer to this age problem Seven years ago, Rolex was six times old as Rose, one year from now, he will be twice old as Rose. Find their present ages.
Answer by longjonsilver(2297)

Let x = Rolex present age
Let y = Rose present age

(x-7) = 6(y-7)
x-7 = 6y - 42
x-6y = -35

(x+1) = 2(y+1)
x+1 = 2y+2
x-2y = 1

So, we have 2 equations:

x-6y = -35
x-2y = 1

Subtract these, to give -4y = -36 --> y = 9

x-2y = 1
x-2(9) = 1
x-18 = 1
x = 19

So present ages are: Rolex = 19 and Rose = 9

jon.